# Non-resonant relaxation of rotating globular clusters

### v2

@@ 5.2 Non-resonant prediction
The difference is much more striking in the presence of rotation (middle and bottom rows of Fig. 10). In both figures, the orbital reshuffling still operates, and the clusters redistribute their orbits towards smoother distributions of inclinations. Yet, the diffusion structures predicted by the NR theory do not align with the $N$-body ones. Indeed, $N$-body simulations (Fig. 9) present “round” structures as one goes from ${\cos I\!=\!0}$ to ${\cos I\!=\!1}$. In addition, the relaxation rate decreases as one considers orbits within to rotation plane, i.e. as one considers ${\cos I\!\to\!\pm 1}$. This is in sharp contrast with the “straighter” structures predicted by the NR theory (Fig. 10), where the relaxation rate does not vary so much as a function of orbital inclination. This seems to be a robust conclusion, as this mismatch remains as one increases the sampling of the various integrations performed in the NR prediction. A similar conclusion is reached in appendix J when considering relaxation in ${(L,\cos I)}$.
Before concluding, let us elaborate further on the failure of the NR theory to model out-of-plane relaxation. For ${\alpha\!\neq\!0}$, recall that ${\cos I\!=\!\pm 1}$ corresponds to orbits within the rotation plane, while ${\cos I\!=\!0}$ corresponds to orbits perpendicular to the rotation plane. A test star orbiting in this rotating cluster will torque with all the other stars of the cluster. Equivalently, this means that this star will typically be subject to a torque with the total angular momentum of the rotating cluster, ${\langle\bm{L}\rangle\!\propto\!\alpha\,\mathbf{e}_{z}}$, the strength of which depends on the test star’s orbital orientation. This is what we observe in Fig. 9. However, the NR theory only takes into account local deflections, and therefore cannot take these coherent interactions into account. This is likely why we obtain a prediction which only weakly depends on orbital inclination (Fig. 10).
Following the work of Meiron & Kocsis (2019) (see also references therein), we anticipate for VRR to play a key role in these coherent interactions. Indeed, VRR, which is driven by persistent torques between orbital planes, might be needed to explain (i) the faster out-of-plane relaxation (section 3.3); (ii) the discrepancies between the $N$-body measurements and the NR predictions. Investigating quantitatively the efficiency of VRR in these systems will be the topic of future investigations.
@@ 6.1 Conclusion
We subsequently turned our interest on out-of-plane relaxation, i.e. the redistribution of orbital inclinations. We pointed out the similarities and differences between the $N$-body measurements and the NR prediction. In both approaches, we observe a systematic reshuffling of orbital inclinations from overpopulated regions to underpopulated ones. As such, the distribution of inclinations get smoother through time, and this process is accelerated by rotation. However, we noted that for for orbits aligned with the cluster’s rotation plane, the local NR theory does not match with the $N$-body measurements. We argue that this mismatch originates from the fact that the NR formalism is sourced by local deflections, which neglect the coherent torques between the orbital planes. Accounting for the (efficient) contribution of persistent torques requires the use of VRR.
@@ 6.2 Perspectives
Here, the NR formalism was used to describe the very onset of relaxation. Yet, the short duration of our simulations and our limited sample of initial conditions prevented us from fully assessing the asymptotic late time orbital distribution. Naturally, it would be interesting to push the $N$-body integrations further in time. While some studies carried out such long-term $N$-body simulations (Tiongco et al., 2020; Livernois et al., 2022), orbital inclinations were, unfortunately, not their focus. Along the same line, it would also be of interest to integrate the 3D FP equation itself. In isotropic clusters, this process has been conducted, e.g. as demonstrated in the work of Vasiliev (2015). However, the scenario involving anisotropic, rotating systems has yet to be comprehensively investigated.
Ultimately, if rotation is sufficiently large, it may induce a flattening of the cluster. In that case, Stäckel systems (Dejonghe & de Zeeuw, 1988) could be used to model realistic flattened rotating structures. In particular, this would still ensure the integrability of the mean potential, i.e. the existence of explicit angle-action coordinates. Tailoring secular theory to such setups will be the topic of future works.
@@ A.2 Arguments of the background distribution
The local in-plane diffusion coefficients, i.e. diffusion coefficients in $E$ and $L$, are given in appendix C of Bar-Or & Alexander (2016). We reproduce them here for completeness. They read
@@ C.1 Angular orbit average
We start from the generic formula of equation (4). There, the average w.r.t. $\varphi$ must be performed with care since $L_{z}^{\prime}$ depends on $\varphi$ (equation 43). Fortunately, taking advantage of the particular structure of the diffusion coefficients in $E$, $L$ and $L_{z}$ (see equations 45 and 49), we only need to compute the following four integrals over $\varphi$
@@
In equation (3), the only dependence w.r.t. $\varphi$ is in the dependence w.r.t. $L_{z}^{\prime}$ in $F_{\mathrm{rot}}$. In the particular case of the LBD (equation 5), we can perform explicitly the average over $\varphi$. Computing the integrals (50) only requires the evaluation of the two non-trivial integrals
@@ Appendix E $N$-body simulations
The simulations presented throughout the main text are performed using the direct $N$-body code NBODY6++GPU (Wang et al., 2015), version 4.1. We use the exact same run parameters as in appendix G of T22, and draw the initial conditions using PlummerPlus.py (Breen et al., 2017). We also refer to the aforementioned appendix for the $N$-body measurements, and to Table 2 for all our binning parameters.
@@ Appendix F Sphericity
We generically define the cluster’s sphericity via ${h\!=\!\lambda_{\min}/\lambda_{\max}}$, which we estimate from $N$-body simulations. To reduce shot noise in that measurement, we average $h$ over realisations as follows. First, we compute the elementary symmetric polynomials ${\alpha\!=\!\lambda_{1}\!+\!\lambda_{2}\!+\!\lambda_{3}}$, ${\beta\!=\!\lambda_{1}\lambda_{2}\!+\!\lambda_{1}\lambda_{3}\!+\!\lambda_{2}\lambda_{3}}$ and ${\gamma\!=\!\lambda_{1}\lambda_{2}\lambda_{3}}$ for every cluster and every timestep. We then average the values of ${(\alpha,\beta,\gamma)}$ over all realisations. From these values, we estimate the eigenvalues $\lambda_{i}$ as the three (positive) roots of the polynomial ${\lambda^{3}\!-\!\langle\alpha\rangle\lambda^{2}\!+\!\langle\beta\rangle\lambda\!-\!\langle\gamma\rangle}$. [^4] The positivity of $\alpha,\beta,\gamma$ – and hence of their averages – ensures the positivity of these roots.
In Fig. 2, we illustrate the evolution of the sphericity, $h$, for various amounts of anisotropy and rotation. We typically find ${h\!\simeq\!0.996}$, i.e. clusters remain reasonably spherically symmetric throughout their evolution. As as visual check of this conclusion, we represent in Fig. 13 the late time stellar distribution of a rotating isotropic globular cluster with ${N\!=\!10^{4}}$. Even close to core collapse, the cluster remains spherically symmetric.
@@ Appendix G Rotating King model
In practice, it is convenient to study long-term relaxation through two-dimensional projections of action space. To do so, we must integrate over one coordinate the 3D FP equation expressed either in ${(J_{\mathrm{r}},L,L_{z})}$ (section D.2) or in ${(J_{\mathrm{r}},L,\cos I)}$ (section D.3). In this appendix, we compute the 2D FP equations in ${(J_{\mathrm{r}},L)}$, ${(J_{\mathrm{r}},\cos I)}$ and ${(L,\cos I)}$ that respectively drive the evolution of the DFs
@@ Appendix I Evaluating the discontinuity at $\cos I=0$
In this appendix, we highlight the sharp discontinuity of the relaxation rate which occurs at $\cos I\!=\!0$ as a result of the discontinuous LBD distribution of orientations (equation 5). We removed from Figures 9, 10, 15 and 16 (see appendices H.2 and H.3). Because equation (5) is discontinuous, we expect that taking the first and second derivatives w.r.t. $\cos I$ will yield a theoretical $\delta^{\prime}(\cos I)$ behaviour around $\cos I\!=\!0$. To that aim, we show in Figure 14 the quantities $D_{\cos\!I}F_{\mathrm{rot}}$ and $D_{\cos\!I\cos\!I}F_{\mathrm{rot}}$, respectively with their first and second w.r.t. $\cos I$.
Figure 14: Representation for a Plummer cluster $(q,\alpha)\!=\!(0,0.25)$ of the $\cos I$ terms, $D_{\cos\!I}F_{\mathrm{rot}}$ (on the top left) and $D_{\cos\!I\cos\!I}F_{\mathrm{rot}}$ (on the top right), respectively with the first (on the bottom left) and second (on the bottom right) derivative w.r.t. $\cos I$. In particular, the bottom right panel displays an approximate $\delta^{\prime}(\cos I)$ behaviour – the amplitude of the jumps tends to infinity as one reduces the step in the finite differentiation scheme – which is the expected theoretical relaxation induced by the $\mathrm{sgn}$ discontinuity of $F_{\mathrm{rot}}$ at $t=0$.
@@
To probe the possible presence of discontinuities and singularities, let us first compute the $\cos I$ component of the 3D flux in ${(J_{\mathrm{r}},L,\cos I)}$, as introduced in section D.3. The dependence of this flux w.r.t. $a$ is illustrated in Fig. 18.
Figure 18: Diffusion flux along $\cos I$ of the 3D FP equation in ${(J_{\mathrm{r}},L,\cos I)}$, for ${J_{\mathrm{r}}\!=\!0.1}$ and ${\cos I\!=\!0.2}$, as a function of $L$ and for various smooth functions $g_{a}$ (equation 83). Here, we consider an isotropic cluster with rotation parameter ${\alpha\!=\!0.25}$. For any smooth $g_{a}$, the flux diverges like ${1/L}$ for ${L\!\to\!0}$. As ${a\!\to\!\infty}$, i.e. as $g_{a}$ tends to the $\mathrm{sgn}$ function, the flux converges to the LBD flux pointwise, which exhibits no divergence.
@@
To estimate the impact of the discontinuity on relaxation, we use $N$-body simulations in Fig. 20 to compare the early relaxation of the discontinuous LBD distribution (equation 5) with its smooth approximation (equation 83 with ${a\!=\!10}$).

### v3

@@ 5.2 Non-resonant prediction
The difference is much more striking in the presence of rotation (middle and bottom rows of Fig. 10). In both figures, the orbital reshuffling still operates, and the clusters redistribute their orbits towards smoother distributions of inclinations. Yet, the diffusion structures predicted by the NR theory do not align with the $N$-body ones. Indeed, $N$-body simulations (Fig. 9) present “round” structures as one goes from ${\cos I\!=\!0}$ to ${\cos I\!=\!1}$. In addition, the relaxation rate decreases as one considers orbits within to rotation plane, that is, as one considers ${\cos I\!\to\!\pm 1}$. This is in sharp contrast with the “straighter” structures predicted by the NR theory (Fig. 10). Nevertheless, one should keep in mind that Fig. 10 cuts out the singularity at ${\cos I\!=\!0}$. The comparison to the $N$-body measurements near ${\cos I\!=\!0}$ is made difficult by the fact that diffusion at early time is very fast in that region, as shown by Fig. 20. As a consequence, the measured relaxation rate may suffer from truncation errors – due to, for example, the finite differentiation scheme – the closer one gets to ${\cos I\!=\!0}$. However, as Fig. 20 shows, the initial discontinuity appears to have little impact on diffusion both far away from ${\cos I\!=\!0}$ and beyond ${t\!=\!0.1\,t_{\mathrm{rh}}}$. The obtention of a more robust measure would require both a finer time difference and a finer binning of action space. These two improvements would require an ensemble-average over much more realisations to reduce dispersion. A similar conclusion is reached in appendix J when considering relaxation in ${(L,\cos I)}$.
Before concluding, let us elaborate further on the failure of the NR theory to model out-of-plane relaxation. For ${\alpha\!\neq\!0}$, recall that ${\cos I\!=\!\pm 1}$ corresponds to orbits within the rotation plane, while ${\cos I\!=\!0}$ corresponds to orbits perpendicular to the rotation plane. A test star orbiting in this rotating cluster will torque with all the other stars of the cluster. Equivalently, this means that this star will typically be subject to a torque with the total angular momentum of the rotating cluster, ${\langle\bm{L}\rangle\!\propto\!\alpha\,\mathbf{e}_{z}}$, the strength of which depends on the test star’s orbital orientation. This is what we observe in Fig. 9. However, the NR theory only takes into account local deflections, and therefore cannot take these coherent interactions into account.
Following the work of Meiron & Kocsis (2019) (see also references therein), we anticipate for VRR to play a key role in these coherent interactions – even after factoring the truncation errors due to the finite difference scheme and the singularity at ${\cos I\!=\!0}$, both discussed earlier in this section. Indeed, VRR, which is driven by persistent torques between orbital planes, might be needed to explain (i) the faster out-of-plane relaxation (section 3.3); (ii) the discrepancies between the $N$-body measurements and the NR predictions. Investigating quantitatively the efficiency of VRR in these systems will be the topic of future investigations.
@@ 6.1 Conclusion
We subsequently turned our interest on out-of-plane relaxation, that is, the redistribution of orbital inclinations. We pointed out the similarities and differences between the $N$-body measurements and the NR prediction. In both approaches, we observe a systematic reshuffling of orbital inclinations from overpopulated regions to underpopulated ones. As such, the distribution of inclinations get smoother through time, and this process is accelerated by rotation. On the one hand, we observed that the initial discontinuity at ${\cos I\!=\!0}$ induced a fast relaxation at small times. Surely, this strongly affects any measurement of the relaxation rate in that region. On the other hand, for orbits aligned with the cluster’s rotation plane – where diffusion is much slower and measurements are much easier – the local NR theory does not match with the $N$-body measurements. We argue that this mismatch originates from the fact that the NR formalism is sourced by local deflections, which neglect the coherent torques between the orbital planes. Accounting for the (efficient) contribution of persistent torques requires the use of VRR.
@@ 6.2 Perspectives
Here, the NR formalism was used to describe the very onset of relaxation. Yet, the short duration of our simulations and our limited sample of initial conditions prevented us from fully assessing the asymptotic late time orbital distribution. Naturally, it would be interesting to push the $N$-body integrations further in time. While some studies carried out such long-term $N$-body simulations (Tiongco et al., 2020; Livernois et al., 2022), orbital inclinations were, unfortunately, not their focus. Along the same line, it would also be of interest to integrate the 3D FP equation itself. In isotropic clusters, this process has been conducted, as demonstrated for instance in the work of Vasiliev (2015). However, the scenario involving anisotropic, rotating systems has yet to be comprehensively investigated.
Ultimately, if rotation is sufficiently large, it may induce a flattening of the cluster. In that case, Stäckel systems (Dejonghe & de Zeeuw, 1988) could be used to model realistic flattened rotating structures. In particular, this would still ensure the integrability of the mean potential, that is the existence of explicit angle-action coordinates. Tailoring secular theory to such setups will be the topic of future works.
@@ Acknowledgements
Finally, we thank our referee, Prof. Douglas Heggie, for his remarks, which contributed greatly to the improvement of this paper.
@@ A.2 Arguments of the background distribution
The local in-plane diffusion coefficients, that is, the diffusion coefficients in $E$ and $L$, are given in appendix C of Bar-Or & Alexander (2016). We reproduce them here for completeness. They read
@@ C.1 Angular orbit average
We start from the generic formula of equation (4). There, the average with respect to $\varphi$ must be performed with care since $L_{z}^{\prime}$ depends on $\varphi$ (equation 43). Fortunately, taking advantage of the particular structure of the diffusion coefficients in $E$, $L$ and $L_{z}$ (see equations 45 and 49), we only need to compute the following four integrals over $\varphi$
@@
In equation (3), the only dependence with respect to $\varphi$ is in the dependence with respect to $L_{z}^{\prime}$ in $F_{\mathrm{rot}}$. In the particular case of the LBD (equation 5), we can perform explicitly the average over $\varphi$. Computing the integrals (50) only requires the evaluation of the two non-trivial integrals
@@ Appendix E $N$-body simulations
The simulations presented throughout the main text were performed using the direct $N$-body code NBODY6++GPU (Wang et al., 2015), version 4.1. We used the exact same run parameters as in appendix G of T22, and drew the initial conditions using PlummerPlus.py (Breen et al., 2017). We also refer to the aforementioned appendix for the $N$-body measurements, and to Table 2 for all our binning parameters.
@@ Appendix F Sphericity
We generically define the cluster’s sphericity via ${h\!=\!\lambda_{\min}/\lambda_{\max}}$, which we estimated from $N$-body simulations. To reduce shot noise in that measurement, we averaged $h$ over realisations as follows. First, we computed the elementary symmetric polynomials ${\alpha\!=\!\lambda_{1}\!+\!\lambda_{2}\!+\!\lambda_{3}}$, ${\beta\!=\!\lambda_{1}\lambda_{2}\!+\!\lambda_{1}\lambda_{3}\!+\!\lambda_{2}\lambda_{3}}$ and ${\gamma\!=\!\lambda_{1}\lambda_{2}\lambda_{3}}$ for every cluster and every timestep. We then averaged the values of ${(\alpha,\beta,\gamma)}$ over all realisations. From these values, we estimated the eigenvalues $\lambda_{i}$ as the three (positive) roots of the polynomial ${\lambda^{3}\!-\!\langle\alpha\rangle\lambda^{2}\!+\!\langle\beta\rangle\lambda\!-\!\langle\gamma\rangle}$. [^4] The positivity of $\alpha,\beta,\gamma$ – and hence of their averages – ensures the positivity of these roots.
In Fig. 2, we illustrate the evolution of the sphericity, $h$, for various amounts of anisotropy and rotation. We typically find ${h\!\simeq\!0.996}$, that is, clusters remain reasonably spherically symmetric throughout their evolution. As as visual check of this conclusion, we represent in Fig. 13 the late time stellar distribution of a rotating isotropic globular cluster with ${N\!=\!10^{4}}$. Even close to core collapse, the cluster remains spherically symmetric.
@@ Appendix G Rotating King model
In practice, it is convenient to study long-term relaxation through two-dimensional projections of action space. To do so, we must integrate over one coordinate the 3D FP equation expressed either in ${(J_{\mathrm{r}},L,L_{z})}$ (appendix D.2) or in ${(J_{\mathrm{r}},L,\cos I)}$ (appendix D.3). In this appendix, we compute the 2D FP equations in ${(J_{\mathrm{r}},L)}$, ${(J_{\mathrm{r}},\cos I)}$ and ${(L,\cos I)}$ that respectively drive the evolution of the DFs
@@ Appendix I Evaluating the discontinuity at $\cos I=0$
In this appendix, we highlight the sharp discontinuity of the relaxation rate that occurs at $\cos I\!=\!0$ as a result of the discontinuous LBD distribution of orientations (equation 5). This singularity has been removed from Figures 10 and 16 (see appendices H.2 and H.3). Because equation (5) is discontinuous, we expect that taking the first and second derivatives with respect to $\cos I$ will yield a theoretical $\delta^{\prime}(\cos I)$ behaviour around $\cos I\!=\!0$. To that aim, we show in Figure 14 the quantities $D_{\cos\!I}F_{\mathrm{rot}}$ and $D_{\cos\!I\cos\!I}F_{\mathrm{rot}}$, respectively with their first and second derivatives with respect to $\cos I$.
Figure 14: Representation for a Plummer cluster $(q,\alpha)\!=\!(0,0.25)$ of the $\cos I$ terms, $D_{\cos\!I}F_{\mathrm{rot}}$ (on the top left) and $D_{\cos\!I\cos\!I}F_{\mathrm{rot}}$ (on the top right), respectively with the first (on the bottom left) and second (on the bottom right) derivative with respect to $\cos I$. In particular, the bottom right panel displays an approximate $\delta^{\prime}(\cos I)$ behaviour – the amplitude of the jumps tends to infinity as one reduces the step in the finite differentiation scheme – which is the expected theoretical relaxation induced by the $\mathrm{sgn}$ discontinuity of $F_{\mathrm{rot}}$ at $t=0$.
@@
To probe the possible presence of discontinuities and singularities, let us first compute the $\cos I$ component of the 3D flux in ${(J_{\mathrm{r}},L,\cos I)}$, as introduced in appendix D.3. The dependence of this flux with respect to $a$ is illustrated in Fig. 18.
Figure 18: Diffusion flux along $\cos I$ of the 3D FP equation in ${(J_{\mathrm{r}},L,\cos I)}$, for ${J_{\mathrm{r}}\!=\!0.1}$ and ${\cos I\!=\!0.2}$, as a function of $L$ and for various smooth functions $g_{a}$ (equation 83). Here, we consider an isotropic cluster with rotation parameter ${\alpha\!=\!0.25}$. For any smooth $g_{a}$, the flux diverges like ${1/L}$ for ${L\!\to\!0}$. As ${a\!\to\!\infty}$, that is, as $g_{a}$ tends to the $\mathrm{sgn}$ function, the flux converges to the LBD flux pointwise, which exhibits no divergence.
@@
To estimate the impact of the discontinuity on relaxation, we used $N$-body simulations in Fig. 20 to compare the early relaxation of the discontinuous LBD distribution (equation 5) with its smooth approximation (equation 83 with ${a\!=\!10}$).